OpenStudy (anonymous):

If I was simplifying (4x^3 + 3x^2 − 6x) − (10x^3 + 3x^2) would the answer be 14x^3 + 6x^2 - 6x

OpenStudy (amistre64):

becareful of the "-"

OpenStudy (anonymous):

no 4-10 =-6 just check

OpenStudy (amistre64):

it distributes thru the (...) and flips all the signs

Parth (parthkohli):

Distributing a '-' is more like distributing a '-1'. Got it?

Parth (parthkohli):

For example, \(-(2x + 3) \Longrightarrow -1(2x + 3) \Longrightarrow (-1)(2x) + (-1)(3) \) Which gets you \(-2x - 3\).

OpenStudy (anonymous):

so (10x^3 + 3x^2) is now (-10x^3 - 3x^2)?

OpenStudy (amistre64):

correct

Parth (parthkohli):

Exactly!

OpenStudy (anonymous):

okay, is the first part left alone?

Parth (parthkohli):

Yep.

OpenStudy (amistre64):

it helps me to stack things up (4x^3 + 3x^2 − 6x) − (10x^3 + 3x^2) 4x^3 + 3x^2 − 6x -10x^3 - 3x^2 -------------------- then just combine

Parth (parthkohli):

It has no '-' sign. It instead has a + sign, which means distributing a +1. \(+(a -b) = +1(a - b) = +a - b = a-b\)

OpenStudy (anonymous):

would the answer be -6x^3 - x^2 - 6x? or is it a positive x^2

OpenStudy (amistre64):

4x^3 + 3x^2 − 6x -10x^3 - 3x^2 --------------------

OpenStudy (amistre64):

what is 3-3?

Parth (parthkohli):

Remember: \(3x^2 - 3x^2 = 0\)

OpenStudy (anonymous):

so would it just be eliminated from the problem or just be like x or something?

Parth (parthkohli):

That comes right from \[3x^2 - 3x^2 \rightarrow (3 - 3)x^2 \rightarrow 0x^2 = 0 \]

Parth (parthkohli):

Yeah, it'd just get removed like breeze.

OpenStudy (amistre64):

0x^2 = 0 by the rule that says when we multiply by zero we end up with zero

OpenStudy (anonymous):

Okay! I understand, thank you guys so much it means a lot!

OpenStudy (amistre64):

good luck :)

Parth (parthkohli):

You're welcome :) I was just a sidekick for amistre heh

OpenStudy (amistre64):

Parth is my conscience, when i go mathically astray he laughs at me lol

Parth (parthkohli):

I need a dictionary for that at the moment. Be right back!

OpenStudy (amistre64):

go - it is the present tense of went ;)

Parth (parthkohli):

lol :)